OpenStudy (kuoministers):

can x^3 + 8 be factorised? if so it is using a3 + b3 = (a + b)(a2 – ab + b2)?

OpenStudy (anonymous):

(x-2)(x^2 + 2x + 4)

OpenStudy (kuoministers):

But the question is x^3 + 8 not minus?

hartnn (hartnn):

yes, and your are correct,a=x and b=2

hartnn (hartnn):

*you

OpenStudy (kuoministers):

Well i was looking to find it in factored from not the answer

OpenStudy (kuoministers):

form*

hartnn (hartnn):

yeah, i didn't give any answer... factored form will be ... ?

hartnn (hartnn):

x^3+2^3 = (...+...)(....-....+...)

OpenStudy (hba):

Well the formula are as follows, \[a^3 + b^3 = (a + b)(a^2 – ab + b^2)\] \[a^3 - b^3 = (a - b)(a^2 + ab + b^2)\]

OpenStudy (kuoministers):

(x+2)(x^2- 2 + 4) Is this right?

hartnn (hartnn):

the 2nd term in 2nd bracket is incorrect.

OpenStudy (kuoministers):

i mean 2x opps (x+2)(x^2- 2x + 4)

hartnn (hartnn):

yeah, now correct....

OpenStudy (hba):

Yeah right

OpenStudy (kuoministers):

thanks

OpenStudy (hba):

Your'e welcome.